A 6.2 kg bowling ball is thrown straight down a lane at 8.4 m/s and makes a head on collision with a 1.6 kg bowling pin initially at rest after the collision the bowling ball is moving in the same direction but has a velocity of only 6.7 m/s what is the speed of the bowling pin after the collision
1
Expert's answer
2020-11-03T10:31:49-0500

We can find the speed of the bowling pin after the collision from the Law of Conservation of Momentum:


"m_bv_{b,i}+m_pv_{p,i}=m_bv_{b,f}+m_pv_{p,f},"

here, "m_b=6.2\ kg" is the mass of the bowling ball, "m_p=1.6\ kg" is the mass of the bowling pin, "v_{b,i}=8.4\ \dfrac{m}{s}" is the initial speed of the bowling ball before the collision, "v_{p,i}=0\ \dfrac{m}{s}" is the initial speed of the bowling pin before the collision, "v_{b,f}=6.7\ \dfrac{m}{s}" is the final speed of the bowling ball after the collision, "v_{p,f}" is the final speed of the bowling pin after the collision.

Then, from this equation we can calculate "v_{p,f}":


"v_{p,f}=\dfrac{m_bv_{b,i}-m_bv_{b,f}}{m_p},""v_{p,f}=\dfrac{6.2\ kg\cdot 8.4\ \dfrac{m}{s}-6.2\ kg\cdot 6.7\ \dfrac{m}{s}}{1.6\ kg}=6.6\ \dfrac{m}{s}."

Answer:

"v_{p,f}=6.6\ \dfrac{m}{s}."